# Fun-ASR Technical Report

### v3

@@ Abstract
In recent years, automatic speech recognition (ASR) has witnessed transformative advancements driven by three complementary paradigms: data scaling, model size scaling, and deep integration with large language models (LLMs). However, LLMs are prone to hallucination, which can significantly degrade user experience in real-world ASR applications. In this paper, we present Fun-ASR, a large-scale, LLM-based ASR system that synergistically combines massive data, large model capacity, LLM integration, and reinforcement learning to achieve state-of-the-art performance across diverse and complex speech recognition scenarios. Moreover, Fun-ASR is specifically optimized for practical deployment, with enhancements in streaming capability, noise robustness, code-switching, hotword customization, and satisfying other real-world application requirements. Experimental results show that while most LLM-based ASR systems achieve strong performance on open-source benchmarks, they often underperform on real industry evaluation sets. Thanks to production-oriented optimizations, Fun-ASR achieves SOTA performance on real application datasets, demonstrating its effectiveness and robustness in practical settings.
Figure 1: Performance comparison (in Accuracy) between our Fun-ASR (7.7B) and Fun-ASR-nano (0.8B) and top-tier ASR and speech-text multimodal models, including Paraformer-v2, Kimi-Audio (8B), FireRedASR (1.1B), and Seed-ASR ($\sim$12B).
@@ 1 Introduction
Data scaling has proven to be a fundamental driver of ASR improvements. The seminal work on Whisper (Radford et al., 2023) provides empirical evidence that ASR performance exhibits a strong positive correlation with the scale of the training data. Their comprehensive experiments demonstrated that increasing the training data volume from 3K hours to over 680K hours results in more than 20-point drop in the English word error rate (WER). This significant improvement underscores the critical role of data diversity and quantity in ASR development, as larger datasets enable models to capture a more comprehensive representation of linguistic and acoustic variations across different languages, accents, speaking styles, and environmental conditions. The availability of massive, multilingual speech corpora has become a cornerstone of modern ASR development, with the most successful systems now leveraging datasets spanning tens of millions of hours.
@@
The third paradigm, deep integration with LLMs, represents a paradigm shift in the ASR methodology. Rather than treating ASR as a standalone task, this approach leverages the rich linguistic knowledge and contextual understanding of LLMs to enhance speech recognition. Models such as Seed-ASR (Bai et al., 2024) and FireRedASR (Xu et al., 2025) have demonstrated that incorporating LLMs can significantly improve ASR performance, particularly in resolving semantic ambiguities and generating more coherent and contextually appropriate transcriptions. These models effectively bridge the gap between speech and text understanding.
@@ 3.1 Pre-taining Data
The pre-training dataset comprises approximately tens of millions hours of audio data, including both unlabeled audio and labeled audio-text data. The unlabeled audio data span a broad range of real-world scenarios in domains such as artificial intelligence, biotechnology, e-commerce, education, entertainment, finance, and mobility. For labeled data, a comprehensive data processing pipeline is employed, which incorporates voice activity detection (VAD), pseudo-label generation by multiple ASR systems (such as Paraformer-V2 (An et al., 2024b), Whisper, and SenseVoice (An et al., 2024a)), followed by inverse text normalization (ITN). The primary languages in the labeled dataset are Chinese and English.
@@ 3.2 Supervised Fine-tuning Data
The supervised fine-tuning (SFT) data consist of approximately millions of hours of data, including: human-transcribed data, pseudo-labeled data, environmental noise data, TTS generated data using CosyVoice3 (Du et al., 2025), simulated streaming data, noise augmented data, and hotword customized data.
@@ 4.1 Pre-training of Audio Encoder
Stage 1: Self-supervised pre-training via Best-RQ framework with pre-trained text LLM initialization. The first stage leverages self-supervised pre-training through the Best-RQ (BERT-based Speech pre-Training with Random-projection Quantizer) framework (Chiu et al., 2022), a state-of-the-art self-supervised learning method for speech representation learning. Best-RQ operates by masking and reconstructing speech units, using a quantization module to discretize continuous representations. This approach enables the model to learn general-purpose speech representations without requiring any labeled data, making it highly scalable to vast amounts of unlabeled audio data. Notably, a key innovation in our implementation of Best-RQ lies in the initialization strategy. Drawing inspiration from our recent findings (An et al., 2024c), where we demonstrate that layers in pre-trained text LLM can effectively initialize the encoder in ASR systems, we initialize the Best-RQ encoder with weights from layers of a pre-trained text LLM—specifically, Qwen3 model (Yang et al., 2025a). This cross-modal initialization strategy is based on the hypothesis that the deep linguistic and semantic knowledge encoded in the LLM can provide a beneficial inductive bias for learning speech representations. We observe that initializing Best-RQ with a pre-trained text LLM significantly accelerates the training convergence and improves the quality of the learned speech representations compared to random initializations.
Stage 2: Supervised pre-training via attention-based encoder-decoder (AED) framework. The second stage involves supervised pre-training of the audio encoder within a conventional attention-based encoder-decoder architecture. This method follows a well-established and empirically validated training paradigm used in our prior works, specifically the SenseVoice-Large model (An et al., 2024a). In this setup, the encoder is trained end-to-end on large-scale labeled ASR datasets, using standard sequence-to-sequence learning objectives. The primary objective of this supervised pre-training phase is to obtain an encoder that learns rich acoustic and linguistic features from transcribed speech data. Once trained, the encoder of the AED framework is used to initialize the audio encoder in the downstream LLM-ASR system. This initialization provides a strong starting point for subsequent joint training of the audio and language components, reducing the need for extensive low-level feature learning from scratch and thereby accelerating the training convergence.
@@ 4.3 Contextual Supervised Fine-tuning
Step 1: Keyword Extraction: To generate contextual information related to the current conversation content, we first extract keywords from its transcript using Qwen3-32B (Yang et al., 2025b). Keywords typically include entities, professional terms, and specific time periods. They are words that ASR systems often fail to recognize.
@@ 4.4.1 The RL Framework for Large Audio-Language Models
We design FunRL, an efficient reinforcement learning (RL) framework tailored for large audio-language models (LALMs). Different from text LLMs, Fun-ASR, as an LALM, incorporates an audio encoder to convert speech into embeddings—a component that is not natively supported by existing RL frameworks such as Verl (Sheng et al., 2024) or Trl (von Werra et al., 2020). As illustrated in Figure 4(a), FunRL orchestrates the audio encoder, rollout, and policy modules using Ray, enabling them to alternately utilize GPU resources. In the audio encoder inference stage, all input audio clips are batched and processed through a Torch-based encoder. The encoder extracts audio embeddings in parallel and transfers the resulting embeddings from GPU to CPU. Subsequently, the SGLang-based LLM rollout takes control of the GPU to generate multiple hypothesis sequences based on the audio embeddings and the instruction text tokens. Each hypothesis is assigned a reward according to the predefined rules, which will be detailed later. Finally, the FSDP-based LLM policy model uses the audio embeddings and the generated hypotheses to compute output probabilities and performs policy optimization via RL. After each update, the optimized policy is synchronized back to the rollout module, ensuring that the RL process remains on-policy.
@@ 5.3 Multilingual ASR
The availability of training data varies widely across different languages. Resource-rich languages such as Chinese and English have abundant data, whereas languages such as Vietnamese and Thai have comparatively limited resources. The primary Fun-ASR model is a Chinese-English model. To improve multilingual ASR performance, we train an additional multilingual Fun-ASR model to support these languages. The current multilingual Fun-ASR model (denoted by Fun-ASR-ML) supports Chinese, English, Vietnamese, Thai, and Indonesian, with a broader language coverage planned for future releases. During training, we downsample the Chinese and English data used for the Chinese-English Fun-ASR model and upsample the Vietnamese, Thai, and Indonesian data to balance the distribution. In total, this multilingual dataset comprises approximately 500K hours of audio. The training methodology is the same as that used for the Chinese-English Fun-ASR model (Section 4).
@@ 5.4 Code-switching
Recognition of code-switched speech has always been a challenge. To optimize ASR performance on Chinese-English code-switched speech, we synthesize the code-switching training data as follows. Firstly, we collect over 40K English key words or phrases, covering common domains such as technology, education, finance, and sports. Secondly, we use the Qwen3 (Yang et al., 2025a) model to generate Chinese-English code-switched texts related to given key words randomly selected from the pool described above. Thirdly, we use a Text-to-Speech model to synthesize speech data in a variety of voices for the LLM-generated code-switched texts, and obtain the code-switched training data.
@@ 6.1 Evaluation Setting
We evaluate Fun-ASR and Fun-ASR-ML on both open-source ASR benchmark datasets and real-world industry evaluation sets. For the open-source evaluation, we use corresponding test sets of AIShell-1 (Bu et al., 2017), AIShell-2(Du et al., 2018), Librispeech(Panayotov et al., 2015), Fleurs(Conneau et al., 2022), WeNetSpeech(Zhang et al., 2022), Gigaspeech2 (Yang et al., 2025c) data sets. These open-source datasets have been publicly available for a long time, increasing the risk of data leakage into model training sets. To ensure a more reliable and leakage-free evaluation, we collect newly uploaded videos from YouTube and Bilibili posted after June 30th, which are then manually transcribed to form an independent test set. For noise robustness evaluation, we use real-world audio recordings captured in various environments, including canteen, dinner, meeting, office, outdoor, park, shop, street, subway, supermarket, and walk-street. These are further categorized by acoustic conditions and topic to better assess performance under diverse and challenging scenarios.
@@ 6.2.1 Overall Results
Whisper- Seed- Seed- Kimi- Step- FireRed- Fun-ASR Fun- Test set large-v3 ASR ASR* Audio Audio2 ASR -nano ASR AIShell1 4.72 0.68 1.63 0.71 0.63 0.54 1.76 1.22 AIShell2 4.68 2.27 2.76 2.86 2.10 2.58 2.80 2.3 Fleurs-zh 5.18 3.43 3.23 3.11 2.68 4.81 3.47 2.64 Fleurs-en 6.23 9.39 9.39 6.99 3.03 10.79 7.95 5.84 Librispeech-clean 1.86 1.58 2.8 1.32 1.17 1.84 1.75 1.57 Librispeech-other 3.43 2.84 5.69 2.63 2.42 4.52 4.37 3.24 WenetSpeech Meeting 18.39 5.69 7.07 6.24 4.75 4.95 8.78 6.49 WenetSpeech Net 11.89 4.66 4.84 6.45 4.67 4.94 6.28 5.46
Table 1: Evaluation results in terms of WER (%) on open-source datasets. The Seed-ASR* results are evaluated using the official API on the volcengine.
Seed- Whisper- FireRed- Kimi- Paraformer Fun-ASR Fun- Test set ASR large-v3 ASR Audio v2 -nano ASR In-house 7.20 16.58 10.10 9.02 8.11 7.26 6.66 Fairfield 4.59 22.21 7.49 10.95 9.55 5.43 4.66 Home Scenario 8.08 18.17 9.67 23.79 6.87 6.02 5.17 Complex Background 12.90 32.57 15.56 15.56 15.19 17.07 11.29 English General 15.65 18.56 21.62 18.12 19.48 15.87 14.22 Opensouce 3.83 7.05 5.31 3.20 6.23 4.65 3.60 Average 8.71 19.19 11.63 13.54 10.91 9.38 7.60
@@ 6.2.2 Streaming ASR Performance
In order to evaluate the streaming ability of our Fun-ASR model, we evaluate the performance on the same test set used when evaluating the offline speech recognition ability of our Fun-ASR model. Table 3 lists the testing results. When comparing with Seed-ASR  (Bai et al., 2024), our Fun-ASR model have better performance at different test sets and test scenario.
Test set Seed-ASR Fun-ASR Fun-ASR-nano In-house 8.64 7.00 7.97 Fairfield 5.51 5.33 6.92 Home Scenario 9.7 5.33 6.51 Complex Background 15.48 12.50 14.83 English General 18.78 14.74 16.70 OpenSouce Test Sets 3.80 3.60 5.13
@@ 6.2.3 Evaluation on Noise Robustness
Fun-ASR Environment w/o NRT w/ NRT NRT + RL canteen 20.67 20.34 19.88 dinner 14.02 9.88 9.55 meeting 6.45 6.27 6.24 office 15.02 11.58 11.42 outdoor 10.12 9.85 9.58 park 13.67 11.88 11.37 shop 12.22 11.48 11.24 street 12.05 10.58 10.86 subway 14.11 13.31 13.29 supermarket 14.27 8.81 8.75 walkstreet 13.89 13.87 13.94 Average 13.32 11.58 11.45

### v4

@@ Abstract
In recent years, automatic speech recognition (ASR) has witnessed transformative advancements driven by three complementary paradigms: data scaling, model size scaling, and deep integration with large language models (LLMs). However, LLMs are prone to hallucination, which can significantly degrade user experience in real-world ASR applications. In this paper, we present Fun-ASR, a large-scale, LLM-based ASR system that synergistically combines massive data, large model capacity, LLM integration, and reinforcement learning to achieve state-of-the-art performance across diverse and complex speech recognition scenarios. Moreover, Fun-ASR is specifically optimized for practical deployment, with enhancements in streaming capability, noise robustness, code-switching, hotword customization, and satisfying other real-world application requirements. Experimental results show that while most LLM-based ASR systems achieve strong performance on open-source benchmarks, they often underperform on real industry evaluation sets. Thanks to production-oriented optimizations, Fun-ASR achieves SOTA performance on real application datasets, demonstrating its effectiveness and robustness in practical settings. The code and models are accessible at https://github.com/FunAudioLLM/Fun-ASR
Figure 1: Performance comparison (in Accuracy) between our Fun-ASR (7.7B) and Fun-ASR-nano (0.8B) and top-tier ASR and speech-text multimodal models.
@@ 1 Introduction
Data scaling has proven to be a fundamental driver of ASR improvements. The seminal work on Whisper (DBLP:conf/icml/RadfordKXBMS23) provides empirical evidence that ASR performance exhibits a strong positive correlation with the scale of the training data. Their comprehensive experiments demonstrated that increasing the training data volume from 3K hours to over 680K hours results in more than 20-point drop in the English word error rate (WER). This significant improvement underscores the critical role of data diversity and quantity in ASR development, as larger datasets enable models to capture a more comprehensive representation of linguistic and acoustic variations across different languages, accents, speaking styles, and environmental conditions. The availability of massive, multilingual speech corpora has become a cornerstone of modern ASR development, with the most successful systems now leveraging datasets spanning tens of millions of hours.
@@
The third paradigm, deep integration with LLMs, represents a paradigm shift in the ASR methodology. Rather than treating ASR as a standalone task, this approach leverages the rich linguistic knowledge and contextual understanding of LLMs to enhance speech recognition. Models such as Seed-ASR (DBLP:journals/corr/abs-2407-04675) and FireRedASR (DBLP:journals/corr/abs-2501-14350) have demonstrated that incorporating LLMs can significantly improve ASR performance, particularly in resolving semantic ambiguities and generating more coherent and contextually appropriate transcriptions. These models effectively bridge the gap between speech and text understanding.
@@ 3.1 Pre-taining Data
The pre-training dataset comprises approximately tens of millions hours of audio data, including both unlabeled audio and labeled audio-text data. The unlabeled audio data span a broad range of real-world scenarios in domains such as artificial intelligence, biotechnology, e-commerce, education, entertainment, finance, and mobility. For labeled data, a comprehensive data processing pipeline is employed, which incorporates voice activity detection (VAD), pseudo-label generation by multiple ASR systems (such as Paraformer-V2 (DBLP:journals/corr/abs-2409-17746), Whisper, and SenseVoice (DBLP:journals/corr/abs-2407-04051)), followed by inverse text normalization (ITN). The primary languages in the labeled dataset are Chinese and English.
@@ 3.2 Supervised Fine-tuning Data
The supervised fine-tuning (SFT) data consist of approximately millions of hours of data, including: human-transcribed data, pseudo-labeled data, environmental noise data, TTS generated data using CosyVoice3 (DBLP:journals/corr/abs-2505-17589), simulated streaming data, noise augmented data, and hotword customized data.
@@ 4.1 Pre-training of Audio Encoder
Stage 1: Self-supervised pre-training via Best-RQ framework with pre-trained text LLM initialization. The first stage leverages self-supervised pre-training through the Best-RQ (BERT-based Speech pre-Training with Random-projection Quantizer) framework (DBLP:conf/icml/ChiuQZYW22), a state-of-the-art self-supervised learning method for speech representation learning. Best-RQ operates by masking and reconstructing speech units, using a quantization module to discretize continuous representations. This approach enables the model to learn general-purpose speech representations without requiring any labeled data, making it highly scalable to vast amounts of unlabeled audio data. Notably, a key innovation in our implementation of Best-RQ lies in the initialization strategy. Drawing inspiration from our recent findings (DBLP:journals/corr/abs-2409-17750), where we demonstrate that layers in pre-trained text LLM can effectively initialize the encoder in ASR systems, we initialize the Best-RQ encoder with weights from layers of a pre-trained text LLM—specifically, Qwen3 model (DBLP:journals/corr/abs-2505-09388). This cross-modal initialization strategy is based on the hypothesis that the deep linguistic and semantic knowledge encoded in the LLM can provide a beneficial inductive bias for learning speech representations. We observe that initializing Best-RQ with a pre-trained text LLM significantly accelerates the training convergence and improves the quality of the learned speech representations compared to random initializations.
Stage 2: Supervised pre-training via attention-based encoder-decoder (AED) framework. The second stage involves supervised pre-training of the audio encoder within a conventional attention-based encoder-decoder architecture. This method follows a well-established and empirically validated training paradigm used in our prior works, specifically the SenseVoice-Large model (DBLP:journals/corr/abs-2407-04051). In this setup, the encoder is trained end-to-end on large-scale labeled ASR datasets, using standard sequence-to-sequence learning objectives. The primary objective of this supervised pre-training phase is to obtain an encoder that learns rich acoustic and linguistic features from transcribed speech data. Once trained, the encoder of the AED framework is used to initialize the audio encoder in the downstream LLM-ASR system. This initialization provides a strong starting point for subsequent joint training of the audio and language components, reducing the need for extensive low-level feature learning from scratch and thereby accelerating the training convergence.
@@ 4.3 Contextual Supervised Fine-tuning
Step 1: Keyword Extraction: To generate contextual information related to the current conversation content, we first extract keywords from its transcript using Qwen3-32B (yang2025qwen3). Keywords typically include entities, professional terms, and specific time periods. They are words that ASR systems often fail to recognize.
@@ 4.4.1 The RL Framework for Large Audio-Language Models
We design FunRL, an efficient reinforcement learning (RL) framework tailored for large audio-language models (LALMs). Different from text LLMs, Fun-ASR, as an LALM, incorporates an audio encoder to convert speech into embeddings—a component that is not natively supported by existing RL frameworks such as Verl (sheng2024hybridflow) or Trl (vonwerra2022trl). As illustrated in Figure 4(a), FunRL orchestrates the audio encoder, rollout, and policy modules using Ray, enabling them to alternately utilize GPU resources. In the audio encoder inference stage, all input audio clips are batched and processed through a Torch-based encoder. The encoder extracts audio embeddings in parallel and transfers the resulting embeddings from GPU to CPU. Subsequently, the SGLang-based LLM rollout takes control of the GPU to generate multiple hypothesis sequences based on the audio embeddings and the instruction text tokens. Each hypothesis is assigned a reward according to the predefined rules, which will be detailed later. Finally, the FSDP-based LLM policy model uses the audio embeddings and the generated hypotheses to compute output probabilities and performs policy optimization via RL. After each update, the optimized policy is synchronized back to the rollout module, ensuring that the RL process remains on-policy.
@@ 5.3 Multilingual ASR
The availability of training data varies widely across different languages. Resource-rich languages such as Chinese and English have abundant data, whereas languages such as Vietnamese and Thai have comparatively limited resources. The primary Fun-ASR model is a Chinese-English model. To improve multilingual ASR performance, we train an additional multilingual Fun-ASR model to support these languages. The previous multilingual Fun-ASR model (denoted by Fun-ASR-ML) supports Chinese, English, Vietnamese, Thai, and Indonesian. For the latest release multilingual model Fun-ASR-ML can support 31 languages, and it also support streaming speech recognition for some languages. In order to support industrial speech recognition application, we also support context bias enhanced speech recognition by using prompt. During training, we downsample the Chinese and English data used for the Chinese-English Fun-ASR model and upsample other languages data in order to balance the distribution. In total, this multilingual dataset comprises approximately 500k hours of audio. The training methodology is the same as that used for the Chinese-English Fun-ASR model (Section 4).
@@ 5.4 Code-switching
Recognition of code-switched speech has always been a challenge. To optimize ASR performance on Chinese-English code-switched speech, we synthesize the code-switching training data as follows. Firstly, we collect over 40K English key words or phrases, covering common domains such as technology, education, finance, and sports. Secondly, we use the Qwen3 (DBLP:journals/corr/abs-2505-09388) model to generate Chinese-English code-switched texts related to given key words randomly selected from the pool described above. Thirdly, we use a Text-to-Speech model to synthesize speech data in a variety of voices for the LLM-generated code-switched texts, and obtain the code-switched training data.
@@ 6.1 Evaluation Setting
We evaluate Fun-ASR and Fun-ASR-ML on both open-source ASR benchmark datasets and real-world industry evaluation sets. For the open-source evaluation, we use corresponding test sets of AIShell-1 (bu2017aishell1opensourcemandarinspeech), AIShell-2(du2018aishell2transformingmandarinasr), Librispeech(7178964), Fleurs(conneau2022fleursfewshotlearningevaluation), WeNetSpeech(zhang2022wenetspeech10000hoursmultidomain), Gigaspeech2 (yang2025gigaspeech2evolvinglargescale) data sets. These open-source datasets have been publicly available for a long time, increasing the risk of data leakage into model training sets. To ensure a more reliable and leakage-free evaluation, we collect newly uploaded videos from YouTube and Bilibili posted after June 30th, which are then manually transcribed to form an independent test set. For noise robustness evaluation, we use real-world audio recordings captured in various environments, including canteen, dinner, meeting, office, outdoor, park, shop, street, subway, supermarket, and walk-street. These are further categorized by acoustic conditions and topic to better assess performance under diverse and challenging scenarios.
@@ 6.2.1 Overall Results
GLM-ASR GLM-ASR Whisper- Seed- Seed- Kimi- Step- FireRed- Fun-ASR Fun- Test set -nano -nano* large-v3 ASR ASR* Audio Audio2 ASR -nano ASR Model Size 1.5B 1.5B 1.6B - - - - 1.1B 0.8B 7.7B OpenSource AIShell1 1.81 2.17 4.72 0.68 1.63 0.71 0.63 0.54 1.80 1.22 AIShell2 - 3.47 4.68 2.27 2.76 2.86 2.10 2.58 2.75 2.39 Fleurs-zh - 3.65 5.18 3.43 3.23 3.11 2.68 4.81 2.56 2.53 Fleurs-en 5.78 6.95 6.23 9.39 9.39 6.99 3.03 10.79 5.96 4.74 Librispeech-clean 2.00 2.17 1.86 1.58 2.8 1.32 1.17 1.84 1.76 1.51 Librispeech-other 4.19 4.43 3.43 2.84 5.69 2.63 2.42 4.52 4.33 3.03 WenetSpeech Meeting 6.73 8.21 18.39 5.69 7.07 6.24 4.75 4.95 6.60 6.17 WenetSpeech Net - 6.33 11.89 4.66 4.84 6.45 4.67 4.94 6.01 5.46
Table 1: Evaluation results in terms of WER (%) on open-source datasets. The Seed-ASR* results are evaluated using the official API on the volcengine and the GLM-ASR-nano* results are evaluated using the opensource checkpoint.
GLM-ASR Whisper-v3 Seed- FireRed- Kimi- Paraformer Fun-ASR Fun- Test set -Nano -large ASR ASR Audio v2 -nano ASR Model Size 1.5B 1.6B - 1.1B 8B 0.2B 0.8B 7.7B OpenSource Nearfield 16.95 16.58 7.20 10.10 9.02 8.11 7.79 6.31 Farfield 9.44 22.21 4.59 7.49 10.95 9.55 5.79 4.34 Complex Background 23.79 32.57 12.90 15.56 15.56 15.19 14.59 11.45 English General 16.47 18.56 15.65 21.62 18.12 19.48 15.28 13.73 Opensource 4.67 7.05 3.83 5.31 3.79 6.23 4.22 3.38 Dialect 54.21 66.14 29.45 52.82 71.94 41.16 28.18 15.21 Accent 19.78 36.03 10.23 14.05 27.20 17.80 12.90 10.31 Lyrics 46.56 54.82 30.26 42.87 65.18 50.14 30.85 21.00 Hiphop 43.32 46.56 29.46 33.88 57.25 43.79 30.87 28.58 Average 26.13 33.39 15.95 22.63 31.00 23.49 16.72 12.70
@@ 6.2.2 Streaming ASR Performance
In order to evaluate the streaming ability of our Fun-ASR model, we evaluate the performance on the same test set used when evaluating the offline speech recognition ability of our Fun-ASR model. Table 3 lists the testing results. When comparing with Seed-ASR  (DBLP:journals/corr/abs-2407-04675), our Fun-ASR model have better performance at different test sets and test scenario.
Test set Seed-ASR Fun-ASR-nano Fun-ASR Nearfield 8.64 8.10 6.75 Farfield 5.51 6.38 4.72 Complex Background 15.48 15.52 12.49 English General 18.78 16.46 14.68 OpenSource 3.80 5.06 4.08 Dialect - 30.72 18.25 Accent - 15.42 11.49 Lyrics - 31.54 22.05 Hiphop - 36.55 28.90
@@ 6.2.3 Evaluation on Noise Robustness
Fun-ASR Environment w/o NRT w/ NRT NRT + RL canteen 20.67 19.66 19.28 dinner 14.02 9.02 8.77 meeting 6.45 6.25 6.05 office 15.02 10.52 10.42 outdoor 10.12 9.63 9.37 park 13.67 11.04 10.89 shop 12.22 10.57 10.46 street 12.05 10.1 9.85 subway 14.11 11.89 11.84 supermarket 14.27 8.03 7.74 walkstreet 13.89 13.34 12.94 Average 13.32 10.91 10.69